That was likely me. I had recentered the beam on the PD I'm using for the armloss measurements, and I probably moved the wrong steering mirror. The transmission from MC2 is sent to a steering mirror that directs it to the MC2 transmission QPD; the transmission from this steering mirror I direct to the armloss MC QPD (the second is what I was trying to adjust).
Note: The MC2 trans QPD goes out to a cable that is labelled MC2 op lev. This confusion should be fixed.
I realigned the MC and recentered the beam on the QPD. Indeed the beam on MC2 QPD was up and left, and the lock was lost pretty quickly, possibly because the beam wasn't centered. Lock was unstable for a while, and I rebooted C1PSL once during this process because the slow machine was unresponsive.
When tweaking the alignment near MC2, take care not to bump the table, as this also chang es the MC2 alignment.
Once the MC was stably locked, I was able to maximize MC transmission at ~15,400 counts. I then centered the spot on the MC2 trans QPD, and transmission dropped to ~14800 counts. After tweaking the alignment again, it was recovered to ~15,000 counts. Gautam then engaged the WFS servo and the beam was centered on MC2 trans QPD, transmission level dropped to ~14,900.
I tried to plot a long trend MC Transmitted today. I could not get farther than 2017 Aug 4
The mode cleaner was misaligned probably due to the earthquake (the drop in the MC transmitted value slightly after utc 7:38:52 as seen in the second plot). The plots show PMC transmitted and MC sum signals from 10th june 07:10:08 UTC over a duration of 17 hrs. The PMC was realigned at about 4-4:15 pm today by rana. This can be seen in the first plot.
The IMC has been misbehaving for the last 5 hours. Why? I turned the WFS servos off. afaik, aaron was the last person to work on the IFO, so i'm not taking any further debugging steps so as to not disturb his setup.
I'm checking out the data this morning, running armloss_AS_calc.py using the parameters Yuki used here.
I made the following changes to scripts (measurement script and calculator script)
I repeated the 'dark' measurements, because I need 20 files to run the script and the measurements before had the window on the scope set larger than the integration time in the script, so it was padded with bad values that were influencing the calculation.
On running the script again, I'm getting negative values for the loss. I removed the beamstops from the PDs, and re-centered the beams on the PDs to repeat the YARM measurements.
The Contec test board with Dsub37Fs was on the top shelf of E7
New all organic machine.
We have no coffee machine.
We are dreaming about it
We still do not have it.
After running this script Friday night, i noticed Saturday that the data hadn't saved. Scrolling up inthe terminal, I couldn't see where I'd run the script, so I thought I'd forgotten to run it as I was making last minute changes to the scope settings Friday before leaving.
Monday it turns out I hadn't forgotten to run the script, but the script itself was getting hung up as it waited for ASS to settle, due to the offset on the ETM PIT or YAW setpoints. The script was waiting until both pitch and yaw settled to below 0.7, but yaw was reading ~15; I think this is normal, and it looks like Yuki had solved this problem by waiting for the DEMOD-OFFSET to become small, rather than just the DEMOD signal to be small. Since this is a solved problem, I think I might be using an old script, but I'm pretty sure I'm running the one in Johannes' folder that Yuki is referencing for example here. The scripts in /yutaro_scripts/ have this DEMOD-OFFSET functionality commented out, and anyway those scripts seem to do the 2D loss maps rather than 1D loss measurements.
In the meantime I blocked the beams and ran the script in DARK mode. The script is saving data in /armloss/data/run_20181105/, and runs with no exceptions thrown.
However, when I try to dither align the YARM, I get an error that "this is not a degree of freedom that has an ASS". I'm alsogetting some exceptions from MEDM about unavailable channels. It must have been something about donatella not initializing, because it's working on pianosa. I turned on YARM ADS from pianosa. Monitoring from dataviewer, I see that LSC-TRY_OUT has some spikes to 0.5, but it's mostly staying near 0. I tried returning to the previous frozen outputs, and also stepping around ETMY-[PIT/YAW] from the IFO_ALIGN screen, but didn't see much change in the behavior of LSC-TRY. I missed the other controls Gautam was using to lock before, and I've also made myself unclear on whether ASS is acting only on angular dof, or also on length.
I unblocked the beams after the DARK run was done.
Some facts which should be considered when doing this measurement and the associated uncertainty:
This result has about 40% of uncertaintities in XARM and 33% in YARM (so big... ).
I'm continuing the arm loss measurements Yuki was making. I'm first familiarizing myself with the procedures for the measurement Johannes describes.
I'm not very familiar with the medm screens, so I'm just kind of poking around and checking with Gautam. I do the following:
I've left the script running.
Let's install Jamie's new Data Viewer
Physical plan is cleaning our roof and gutters today.
It took at least ten years to rust away.
Steve reported to me that the CC1 Hornet gauge was not reporting the IFO pressure after some cable tracing at EX. I found that the power to the unit had been accidentally disconnected. I re-connected the power and manually turned on the HV on the CC gauge (perhaps this can be automated in the new vacuum paradigm). IFO pressure of 8e-6 torr is being reported now.
Chub Osthelder received 40m specific basic safety traning today.
Gautam & Steve,
Our controller is back with Osaka maintenace completed. We swapped it in this morning.
TP-1 Osaka maglev controller [ model TCO10M, ser V3F04J07 ] needs maintenance. Alarm led on indicating that we need Lv2 service.
The turbo and the controller are in good working order.
Our maintenance level 2 service price is $...... It consists of a complete disassembly of the controller for internal cleaning of all ICB’s, replacement of all main board capacitors, replacement of all internal cooling units, ROM battery replacement, re-assembly, and mandatory final testing to make sure it meets our factory specifications. Turnaround time is approximately 3 weeks.
RMA 5686 has been assigned to Caltech’s returning TC010M controller. Attached please find our RMA forms. Complete and return them to us via email, along with your PO, prior to shipping the cont
Osaka Vacuum USA, Inc.
510-770-0100 x 109
our TP-1 TG390MCAB is 9 years old. What is the life expectancy of this turbo?
The Osaka maglev turbopumps are designed with a 100,000 hours(or ~ 10 operating years) life span but as you know most of our end-users are
running their Osaka maglev turbopumps in excess of 10+, 15+ years continuously. The 100,000 hours design value is based upon the AL material being rotated at
the given speed. But the design fudge factor have somehow elongated the practical life span.
We should have the cost of new maglev & controller in next year budget. I put the quote into the wiki.
As a part of the preparation for the replacement of c1susaux with Acromag, I made inspection of the coil-osem transfer function measurements for the vertex SUSs.
The TFs showed typical f^-2 with the whitening on except for ITMY UL (Attachment 1). Gautam told me that this is a known issue for ~5 years.
We made a thorough inspection/replacement of the components and identified the mechanism of the problem.
It turned out that the inputs to MAX333s are as listed below.
The switching voltage for UL is obviously incorrect. We thought this comes from the broken BIO board and thus swapped the corresponding board. But the issue remained. There are 4 BIO boards in total on c1sus, so maybe we have replaced a wrong board?
Initially, we thought that the BIO can't drive the pull-up resistor of 5KOhm from 15V to 0V (=3mA of current). So I have replaced the pull-up resistor to be 30KOhm. But this did not help. These 30Ks are left on the board.
To do for Green Locking in YARM:
The auto-alignment servo should be completed. This servo requires many parameters to be optimized: demodulation frequency, demodulation phase, servo gain (for each M1/2 PIT/YAW), and matrix elements which can remove PIT-YAW coupling.
The main laser went off when PSL doors were opened-closed. It was turned back on and the PSL is locked.
The scripts for measuring armloss are in the directory "/opt/rtcds/caltech/c1/scripts/lossmap_scripts/armloss_scope".
Final Procedure Report for Green Locking in YARM:
The current setup of AUX Y-arm Green locking has to be improved because:
What to do
Steve & Bob,
Bob removed the head cover from the housing to inspect the condition of the the tip seal. The tip seal was fine but the viton cover seal had a bad hump. This misaligned the tip seal and it did not allow it to rotate.
It was repositioned an carefully tithened. It worked. It's starting current transiant measured 28 A and operational mode 3.5 A
This load is normal with an old pump. See the brand new DIP7 drypump as spare was 25 A at start and 3.1 A in operational mode. It is amazing how much punishment a slow blow ceramic 10A fuse can take [ 0215010.HXP ]
In the future one should measure the current pick up [ transient <100ms ] after the the seal change with Fluke 330 Series Current Clamp
It was swapped in and the foreline pressure dropped to 24 mTorr after 4 hours. It is very good. TP3 rotational drive current 0.15 A at 50K rpm 24C
Gautam and Steve,
Our TP3 drypump seal is at 360 mT [0.25A load on small turbo] after one year. We tried to swap in old spare drypump with new tip seal. It was blowing it's fuse, so we could not do it.
Noisy aux drypump turned on and opened to TP3 foreline [ two drypumps are in the foreline now ] The pressure is 48 mT and 0.17A load on small turbo.
but there's one weirdness: It get's the channel offset wrong. However this doesn't matter in our measurement because we're subtracting the dark level, which sees the same (wrong) offset.
When you do this measurement with oscilloscope, take care two things:
I used these values for measuring armloss:
then the uncertainties reported by the individual measurements are on the order of 6 ppm (~6.2 for the XARM, ~6.3 for the YARM). This accounts for fluctuations of the data read from the scope and uncertainties in mode-matching and modulation depths in the EOM. I made histograms for the 20 datapoints taken for each arm: the standard deviation of the spread is over 6ppm. We end up with something like:
XARM: 123 +/- 50 ppm
YARM: 152+/- 50 ppm
In the previous measurement, the fluctuation of each power was 0.1% and the fluctuation of P(Locked)/P(misaligned) was also 0.1%. Then the uncertainty was small. On the other hand in my measurement, the fluctuation of power is about 2% and the fluctuation of P(Locked)/P(misaligned) is 2%. That's why the uncertainty became big.
We want to measure tiny value of loss (~100ppm). So the fluctuation of P(Locked)/P(misaligned) must be smaller than 1.6%.
(Edit on 10/23)
I think the error is dominated by systematic error in scope. The data of beam power had only 3 degits. If P(Locked) and P(misaligned) have 2% error, then
You have to check the configuration of scope.
DASWG is not what we want to use for config; we should use the K. Thorne LLO instructions, like I did for ROSSA.
pianosa has been upgraded to SL7. I've made a controls user account, added it to sudoers, did the network config, and mounted /cvs/cds using /etc/fstab. Other capabilities are being slowly added, but it may be a while before this workstation has all the kinks ironed out. For now, I'm going to follow the instructions on this wiki to try and get the usual LSC stuff working.
the script "armloss_AS_calc.py",
Some changes were made in the script for getting the signals of beam power:
In the yesterday measurement the beam power of ASDC is higher when locked than when misaligned and I wrote it maybe caused by over-coupled cavity. Then I did a calculation as following to explain this:
I ran the script for measuring arm-loss and calculated rough Y-arm round trip loss temporally. The result was 89.6ppm. (The error should be considered later.)
The measurement was done as follows:
('AS_DARK =', '0.0019517200000000003') #dark noise at ASDC
('MC_DARK =', '0.02792') #dark noise at MC2 trans
('AS_LOCKED =', '2.04293') #beam power at ASDC when the cavity was locked
('MC_LOCKED =', '2.6951620000000003')
('AS_MISALIGNED =', '2.0445439999999997') #beam power at ASDC when the cavity was misaligned
('MC_MISALIGNED =', '2.665312')
#normalized beam power
We want to measure the pressure gradient in the 40m IFO
Our old MKS cold cathodes are out of order. The existing working gauge at the pumpspool is InstruTech CCM501
The plan is to purchase 3 new gauges for ETMY, BS and MC2 location.
Basic cold cathode or Bayard-Alpert Pirani
With Gautam's help, Y-arm was locked. Then I ran the script "armloss_dcrefl_asdcpd_scope.py" which gets the signals from oscilloscope. It ran and got data, but I found some problems.
Anyway, I got the data needed so I will calculate the loss after converting the format.
This is the procedure I follow when I take these measurements for the XARM (symmetric under XARM <-> YARM):
Information for the armloss measurement:
Note: The scripts uses httplib2 module. You have to install it if you don't have.
The locked arms are needed to calculate armloss but the alignment of PMC is deadly bad now. So at first I will make it aligned. (Gautam aligned it and PMC is locked now.)
gautam: The PMC alignment was fine, the problem was that the c1psl slow machine had become unresponsive, which prevented the PMC length servo from functioning correctly. I rebooted the machine and undid the alignment changes Yuki had made on the PSL table.
I connected DAC - AIboard - PZTdriver - PZT mirrors and made sure the PZT mirrors were moving when changing the signal from DAC. Tomorrow I will prepare alignment servo with green beam for Y-arm.
[ Yuki, Gautam, Steve ]
To align the green beam in Y-end these hardware were installed:
I made sure that DAC CH9~16 and cable to AI-board worked correctly.
When we applied +100V to PZT driver and connected DAC, AI-board and PZT drive, the output voltage of the driver was not correct. I'll check it tomorrow.
Setting up c1asy:
Now Yuki can work on copying the simulink model (copy c1asx structure) and implementing the autoalignment servo.
While pointing Yuki to the c1asx servo system, I noticed that the filter file for c1asx is missing in the usual chans directory. Why? Backups for it exist in the filter_archive subdirectory. But there is no current file. Clearly this doesn't seems to affect the realtime code execution as the ASX model seems to run just fine. I copied the latest backup version from the archive area into the chans directory for now.
I measured it with the wrong setting of a powermeter. The correct ones are here:
After installation I measured these power again.
There is a little power loss. That may be because of adding one lens in the beam path. I think it is allowable margin.
I finished installation of optics in the Y-end and recovered green locking. Current ALS-TRY_OUTPUT is about 0.25, which is lower than before. So I still continue the alignment of the beam. The simulation code was attached. (Sorry. The optic shown as QWP2 is NOT QWP. It's HWP.)
To facilitate Yuki's alignment of the EY green beam into the Yarm cavity, I have changed the LSC triggering and PowNorm settings to use only the reflected light from the cavity to do the locking of Arm Cavity length to PSL. Running the configure script should restore the usual TRY triggering settings. Also, the X arm optics were macroscopically misaligned in order to be able to lock in this configuration.
I designed a new layout. It has good mode-matching efficiency, reasonable sensitivity to component positions, good Gouy phase separation. I'm setting optics in the Y-end table. The layout will be optimized again after finishing (rough) installation. (The picture will be posted later)
What about just copying the Xend layout? I think it has good MM (per calculations), reasonable (in)sensitivity to component positions, good Gouy phase separation, and I think it is good to have the same layout at both ends. Since the green waist has the same size and location in the doubling crystal, it should be possible to adapt the X end solution to the Yend table pretty easily I think.
The setup I designed is here. It can bring 100% mode-matching and good separation of degrees of TEM01, however I found a probrem. The picture of setup is attached #3. You can see the reflection angle at Y7 and Y8 is not appropriate. I will consider the schematic again.
The calculated conversion efficiency of SHG crystal is 1.2%W.
The SHG crystal has the conversion efficiency of ~2%W (i.e. if you have 1W input @1064, you get 2% conversion efficiency ->20mW@532nm)
It is not possible to produce 0.58mW@532nm from 20.9mW@1064nm because this is already 2.8% efficiency.
Before changing setup at Y-end table, I measured the status value of the former setup as follows. These values will be compared to those of upgraded setup.
(These numbers are shown in the attachment #2.)
[ Yuki, Gautam ]
I made a cable which connects DAC port (40 pins) and AI board (25 pins). I will check if it works.
Tomorrow I will change setup for improvement of AUX Y-end green locking. Any optics for IR will not be moved in my design, so this work doesn't affect Y-arm locking with main beam.
While doing this work, I will do:
I improved Anti-Imaging board (D000186-Rev.D), which will be put between DAC port and PZT driver board.
It had notches at f = 16.6 kHz and 32.7 kHz, you can see them in the plot attached. So I replaced some resistors as follows:
Then the notch moved to 65.9 kHz (> sampling frequency of DAC = 64 kHz, good!).
(The plot enlarged around the notch frequency and the plot of all channels will be posted later.)
All electronics and optics seem to be ready.
Interim Procedure Report:
We have been working on double checking the noise budget calculations. We wanted to evaluate the amount of squeezing for a few different scenarios that vary in cost and time. Here are the findings:
All calculations done with
Main unbudgeted noises:
Threat matrix has been updated.
I calibrated PZT mirrors. The ROUGH result was attached. (Note that some errors and trivial couplings coming from inclination of QPD were not considered here. This should be revised and posted again.)
The PZT mirrors I calibrated were:
I did the calibration as follows:
The calibration factor was
CVI-pitch: 0.089 mrad/V
CVI-yaw: 0.096 mrad/V
Laseroptic-pitch: 0.062 mrad/V
Laseroptic-yaw: 0.070 mrad/V
Previous calibration of the same mirrors, elog:40/8967
I've been plugging away at Altium prototyping the high-voltage bias idea, this is meant to be a progress update.
I need to get footprints for some of the more uncommon parts (e.g. PA95) from Rich before actually laying this out on a PCB, but in the meantime, I'd like feedback on (but not restricted to) the following:
I also don't have a good idea of what the PCB layer structure (2 layers? 3 layers? or more?) should be for this kind of circuit, I'll try and get some input from Rich.
*Updated with current noise (Attachment #2) at the output for this topology of series resistance of 25 kohm in this path. Modeling was done (in LTspice) with a noiseless 25kohm resistor, and then I included the Johnson noise contribution of the 25k in quadrature. For this choice, we are below 1pA/rtHz from this path in the band we care about. I've also tried to estimate (Attachment #3) the contribution due to (assumed flat in ASD) ripple in the HV power supply (i.e. voltage rails of the PA95) to the output current noise, seems totally negligible for any reasonable power supply spec I've seen, switching or linear.